OpenStudy (anonymous):

if f(x)=2x-3 and g(x)=x , what is f(g(2)) A) -2 B) 0 C) 1 D) 4

hartnn (hartnn):

g(2) can be found by putting x=2 in g(x) so, g(2)=... ?

hartnn (hartnn):

@alanis are you there ?

OpenStudy (anonymous):

yeah im here

hartnn (hartnn):

so, can you find g(2) =... ?

OpenStudy (anonymous):

no im hazy on how to do this

hartnn (hartnn):

you just need to replace 'x' by 2 in g(x)=x to get g(2) its really simple, try it.

OpenStudy (anonymous):

so u just times 2 by 2 and get 4

hartnn (hartnn):

where did 2 times 2 come from ?? g(x) = x when i put x=2, i get g(2) =2 go this ?

hartnn (hartnn):

Note, i have just replaced 'x' by '2'

OpenStudy (anonymous):

ok so i did all the math and i came up with D) 4 is that the correct answer?

hartnn (hartnn):

no, its incorrect. show your math work, i'll correct it.

OpenStudy (anonymous):

\[f(g(2)) g(2)=(2)=2 f(g(2))=4=2(4)-3\]

hartnn (hartnn):

no, thats not correct. g(2)=2 so, f(g(2)) =f(2) = 2(2)-3 = .... ?

OpenStudy (anonymous):

so itll be 1

hartnn (hartnn):

yes!

OpenStudy (anonymous):

thanks

hartnn (hartnn):

welcome ^_^